The length of a rectangle is four less than twice the width. The perimeter is 46 inches. Find the dimensions of the rectangle.
Found 2 solutions by rfer, checkley77:
Answer by rfer(10432)

p=2w+2l
w=x=9
l=2x-4=14
--------------
2x+2(2x-4)=46
2x+4x-8=46
6x=54
x=9
------------
18+28=46

Answer by checkley77(12569)

L=2W-4
P=2L+2W
46=2(2W-4)+2W
46=4W-8+2W
46+8=6W
54=6W
W=54/6
W=9 FOR THE WIDTH.
L=2*9-4
L=18-4
L=14 FOR THE LENGTH.
PROOF:
46=2*14+2*9
46=28+18
46=46